• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Deur

    Log entry time 00:02:36 on November10,2009

    Entry number 298664

    keyword=Swing shift summary

    4:00 Shift started with the Hall in controlled access to fix the left HRS 
    4:00 Technician leaving the Hall (left dipole still down). We asked MCC for the 
    beam back. They said they are doing a quad scan so it will take about 30min 
    before they can put us back to beam permit. 
    4:25 Turned the VDC back on.
    5:25 MCC called. They will establish the beam in the hall.
    5:32 Beam is established. We move to C12 target. Request 20 uA.
    17:42 we have 20 uA in the hall for detector and DAQ check. We start 
    non-production run #4892 (right arm).
    17:48 Detectors & DAQ are fine. We stop the run, and called MCC to mask the 
    target in order to move to empty. Then Tehy will start Compton tune.
    18:25 Jack came and cleared a fault for left arm Q1. We are ramping up Q1 (the 
    point is that we have symmetric fringe field for using the lumi)
    18:28 Compton HV trip (followed by many others)
    20:25. MCC called. Beam is tune on the target. They will send high current beam 
    now, straight (i.e not through the chicane).
    20:35 MCC called and say they are happy with their tune. They are trying now 
    throught the chicane.
    20:55 We called MCC for update. They have back out of the chicane and are 
    redoing the straight-trhough part.
    21:53 MCC called. They are done with the straight tunning. Now they will try 
    through the chicane.
    22:15 Compton tune is done. We move the target from empty to carbon. Starting 
    target edge scans.
    23:55 Target study completed. Preparing for production. Ramping down the VDC HV.
    00:00 requesting 100 uA.